find the value of m and n that make the equation true.
(6-m)+3ni=-12+27i

Answer by stanbon(57377)

find the value of m and n that make the equation true.
(6-m)+3ni=-12+27i
The left side and the right side are complex numbers.
Each number has a Real Number part and aa Imaginary Number part.
Equate the Real Number parts to solve for "m":
6-m = -12
m = 18
---------------
Equate the Imaginary Number parts to solve for "n":
3ni = 27i
n = 9
------------
Cheers,
Stan H.